# Cyclotomic generating functions

### v2

@@ 2.1. Cyclotomic polynomials
Also, $\Phi_{p}(q)=1+q+q^{2}+\cdots+q^{p-1}=[p]_{q}$ if and only if $p$ is prime. Note, $\Phi_{p^{k}}(1)=p$, otherwise if $n>1$ is not a prime power, then $\Phi_{n}(1)=1$. The number of complex roots and therefore the degree of $\Phi_{n}(q)$ is given by Euler’s totient function, $\varphi(n)$, so for $n\geq 3$ the degree of $\Phi_{n}(q)$ is even and bounded by $\sqrt{n/2}\leq\varphi(n)\leq n-1$. For $n\geq 2$, the constant term is $\Phi_{n}(0)=1$, and the coefficients are palindromic in the sense that $\Phi_{n}(q)=q^{\varphi(n)}\Phi_{n}(q^{-1})$.
2.2. Probabilistic generating functions and cumulants
We now briefly review the background from probability related to CGF’s. See [Bil95] for more details or [BKS20a, §2] for a review aimed at a combinatorial audience.
The probability generating function of a discrete random variable $\mathcal{X}$ supported on $\mathbb{Z}_{\geq 0}$ is
@@ 2.2. Probabilistic generating functions and cumulants
If $\mathcal{X}$ is a non-negative-integer-valued statistic on a combinatorial set $S$ which is sampled uniformly, then the probability generating function of $\mathcal{X}$ as a random variable is the same as the ordinary generating function of $\mathcal{X}$, up to a scale factor:
@@
where $\mu_{d}\coloneqq\mathbb{E}[\mathcal{X}]^{d}$ is the $d$th moment. The central moments of $\mathcal{X}$ are $\alpha_{d}\coloneqq\mathbb{E}[(\mathcal{X}-\mu)^{d}]$, where $\mu=\mu_{1}$ is the mean and $\alpha_{2}=\sigma^{2}$ is the variance. The characteristic function of $\mathcal{X}$ is
@@
$f(q)=\frac{\prod_{k=1}^{m}[a_{k}]_{q}}{\prod_{k=1}^{m}[b_{k}]_{q}}=\sum c_{k}q^{k}\in\Phi^{+},$
so in particular each $c_{k}\in\mathbb{Z}_{\geq 0}$. Let $\mathcal{X}$ be a discrete random variable with $\mathbb{P}[\mathcal{X}=k]=c_{k}/f(1)$. Then the $d$th cumulant of $\mathcal{X}$ is
(2.5) $\kappa_{d}^{\mathcal{X}}=\frac{B_{d}}{d}\sum_{k=1}^{m}(a_{k}^{d}-b_{k}^{d}),$
@@
(2.6) $\alpha_{d}=\sum_{\begin{subarray}{c}|\lambda|=d\\ \text{has all parts even}\end{subarray}}\frac{d!}{z_{\lambda}}\,\prod_{i=1}^{\ell(\lambda)}\frac{B_{\lambda_{i}}}{\lambda_{i}!}\left[\sum_{k=1}^{m}\left(a_{k}^{d}-b_{k}^{d}\right)\right],$
@@
(2.7) $\mu_{d}=\sum_{\begin{subarray}{c}|\lambda|=d\\ \text{has all parts either}\\ \text{even or size $1$}\end{subarray}}\frac{d!}{z_{\lambda}}\,\prod_{i=1}^{\ell(\lambda)}\frac{B_{\lambda_{i}}}{\lambda_{i}!}\left[\sum_{k=1}^{m}\left(a_{k}^{d}-b_{k}^{d}\right)\right].$
Definition 2.2.
@@
Observe that $\mathcal{X}^{*}$ has mean $\mu^{*}=0$ and variance ${\sigma^{*}}^{2}=1$. The moments and central moments of $\mathcal{X}^{*}$ agree for $d\geq 2$ and are given by
@@
Similarly, the cumulants of $\mathcal{X}^{*}$ are given by $\kappa_{1}^{*}=0$, $\kappa_{2}^{*}=1$, and $\kappa_{d}^{*}=\kappa_{d}/\sigma^{d}$ for $d\geq 2$.
Definition 2.3.
@@
Theorem 2.4 (Lévy’s Continuity Theorem, [Bil95, Theorem 26.3]).
@@
Theorem 2.5 (Frechét–Shohat Theorem, [Bil95, Theorem 30.2]).
@@
By 2.4, we may test for asymptotic normality by checking if the standardized characteristic functions tend pointwise to the characteristic function of the standard normal. Likewise by 2.5 we may instead perform the check on the level of individual standardized moments, which is often referred to as the method of moments. By the polynomial relationship between moments and cumulants, we may further replace the moment condition (2.8) with the cumulant condition
(2.9) $\lim_{n\to\infty}\kappa_{d}^{\mathcal{X}_{n}}=\kappa_{d}^{\mathcal{X}}.$
@@
Corollary 2.6.
A sequence $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ of real-valued random variables on finite sets is asymptotically normal if for all $d\geq 3$ we have
(2.10) $\lim_{n\to\infty}\frac{\kappa^{\mathcal{X}_{n}}_{d}}{(\sigma^{\mathcal{X}_{n}})^{d}}=0.$
In fact, the converse of the Frechét–Shohat theorem holds for cyclotomic generating functions. See 5.2 below, which builds on [HZ15, Lem. 2.8]. Furthermore, we have the following simplified test for asymptotic normality due to Hwang and Zacharovas.
Theorem 2.7.
@@
(2.11) $\lim_{n\to\infty}\frac{\kappa^{\mathcal{X}_{n}}_{4}}{(\sigma^{\mathcal{X}_{n}})^{4}}=0.$
2.3. Formal cumulants
@@
(2.12) $\log f(e^{t})=\sum_{d=1}^{\infty}\kappa_{d}(f)\frac{t^{d}}{d!}$
@@
(2.13) $\kappa_{d}(f)=\frac{B_{d}}{d}\left(\sum_{k=1}^{m}a_{k}^{d}-\sum_{k=1}^{m}b_{k}^{d}\right).$
@@
(2.14) $\phi_{f}(t)\coloneqq f(e^{it})/f(1)=\frac{\prod_{k=1}^{m}[a_{k}]_{e^{it}}/a_{k}}{\prod_{k=1}^{m}[b_{k}]_{e^{it}}/b_{k}}=e^{-it\mu}\prod_{j=1}^{m}\frac{\operatorname{sinc}(a_{j}t/2)}{\operatorname{sinc}(b_{j}t/2)},$
@@
(2.15) $\kappa_{d}(\Phi_{n}(q))=\frac{B_{d}}{d}\sum_{k\mid n}\mu(n/k)k^{d}=\frac{B_{d}}{d}J_{d}(n).$
@@
(2.16) $\sum_{k\mid n}\mu(n/k)k^{d}=n^{d}\prod_{\begin{subarray}{c}\text{$p$ prime}\\ p\mid n\end{subarray}}\left(1-\frac{1}{p^{d}}\right).$
@@
Corollary 2.9.
[HZ15, Cor. 3.1] Let $f\in\mathbb{R}_{\geq 0}[q]$ be any non-zero polynomial such that all of its complex roots have modulus $1$. The corresponding odd cumulants $\kappa_{2d-1}(f)$ vanish after the first. The corresponding even cumulants $\kappa_{2d}(f)$ alternate in sign according to $(-1)^{d-1}\kappa_{2d}(f)\geq 0$.
2.4. Generalized uniform sum distributions
@@
Definition 2.10.
Let $\mathbf{t}=(t_{1},t_{2},\ldots)$ be a sequence of non-negative real numbers. For $p\in\mathbb{R}_{\geq 1}$, the $p$-norm of $\mathbf{t}$ is $|\mathbf{t}|_{p}\coloneqq\left(\sum_{k=1}^{\infty}t_{k}^{p}\right)^{1/p}$. We also set $|\mathbf{t}|_{\infty}\coloneqq\sup_{k}t_{k}$.
The $p$-norm has many nice properties. It is well-known (e.g. [MV97, Ex. 7.3, p.58]) that if $1\leq p\leq q\leq\infty$, then $|\mathbf{t}|_{p}\geq|\mathbf{t}|_{q}$, and that if $|\mathbf{t}|_{p}<\infty$, then $\lim_{q\to\infty}|\mathbf{t}|_{q}=|\mathbf{t}|_{\infty}$. Thus, if $\mathbf{t}$ is weakly decreasing, $|\mathbf{t}|_{\infty}=\sup_{k}t_{k}=t_{1}$.
The sequence space with finite $p$-norm $\mathbf{\ell}_{p}\coloneqq\{\mathbf{t}=(t_{1},t_{2},\ldots)\in\mathbb{R}_{\geq 0}^{\mathbb{N}}:|\mathbf{t}|_{p}<\infty\}$ is commonly used in functional analysis and statistics. Here we define a related concept for analyzing sums of central continuous uniform random variables.
Definition 2.11.
@@
~$\widetilde{\ell}_{p}\coloneqq\{\mathbf{t}=(t_{1},t_{2},\ldots):t_{1}\geq t_{2}\geq\cdots\geq 0,|\mathbf{t}|_{p}<\infty\}.$
The elements of ~$\widetilde{\ell}_{p}$ may equivalently be thought of as the set of countable multisets of non-negative real numbers with finite $p$-norm. Any finite multiset of non-negative real numbers can be considered as an element of ~$\widetilde{\ell}_{p}$ with finite support by sorting the multiset and appending 00’s. The multisets in ~$\widetilde{\ell}_{p}$ are uniquely determined by their $p$-norms. In fact, any sequence of $p$-norm values injectively determines the multiset provided the sequence goes to infinity.
Definition 2.12.
@@
The uniform sum random variables have nice cumulant formulas which are similar to the CGF distributions. It was shown in [BS22, Lem. 3.11] that for $d\geq 2$ and $\mathbf{t}=(t_{1},t_{2},\ldots)\in\widetilde{\ell}_{2}$, we have
(2.17) $\kappa_{d}^{\mathcal{S}_{\mathbf{t}}}=\frac{B_{d}}{d}\sum_{k=1}^{\infty}(t_{k})^{d}=\frac{B_{d}}{d}|\mathbf{t}|_{d}^{d}.$
Theorem 2.13.
@@
It is not known what are all possible limiting distributions of families of CGF polynomials. By [BS22, Cor. 3.17], we know that all standardized uniform sum distributions $\mathcal{S}_{\mathbf{t}}^{*}$ do occur as limiting distributions coming from the hook length formulas for linear extensions of forests due to Björner and Wachs [BW89]. In fact, all standardized DUSTPAN distributions can occur as limits of CGF distributions. These are distributions of the form $\mathcal{S}_{\mathbf{t}}+\mathcal{N}(0,\sigma^{2})$, assuming the two random variables are independent, $\mathbf{t}\in\widetilde{\ell}_{2}$, and $\sigma^{2}\coloneqq 1-|\mathbf{t}|_{2}^{2}/12\in\mathbb{R}_{\geq 0}$.
3. Algebraic considerations
3.1. Equivalent characterizations of CGF’s
@@
If $f(q)=\prod_{j=1}^{n}(q-z_{j})$ for $z_{1},\ldots,z_{n}\in\mathbb{C}$, define $f_{k}(q)=\prod_{j=1}^{n}(q-z_{j}^{k})$ for each positive integer $k$. The coefficients of $f_{k}(q)$ are elementary symmetric polynomials in the $z_{j}^{k}$ and each $|z_{j}^{k}|\leq 1$ by hypothesis, so the coefficients of $f_{k}(q)$ are bounded in modulus by binomial coefficients. This expresses them as symmetric functions of Galois conjugates of $f$, so they belong to the fixed field of the Galois group, namely the base field $\mathbb{Q}$. Since $f$ is monic, its roots $z_{1},\ldots,z_{n}$ are algebraic integers as are all sums of products of the $z_{i}$’s. Therefore, the coefficients of each $f_{k}(q)$ are also algebraic integers, and since the only algebraic integers in $\mathbb{Q}$ are the integers themselves, we observe $f_{k}(q)\in\mathbb{Z}[q]$. The list $f_{1},f_{2},\ldots$ must thus eventually repeat. We may as well suppose $f=f_{1}$ is repeated, so that taking $k$th powers for some $k>1$ permutes the $z_{i}$’s. Hence, taking $k$th powers $n!$ times implies $z_{i}=z_{i}^{kn!}$ for all $i$, and the result follows. ∎
@@
(3.1) $\Phi_{n}(q)=\prod_{d\mid n}(q^{d}-1)^{\mu(n/d)}=\prod_{d\mid n}[d]_{q}^{\mu(n/d)}=\prod_{j=1}^{m}\frac{1-q^{a_{j}}}{1-q^{b_{j}}}=\prod_{j=1}^{m}\frac{[a_{j}]_{q}}{[b_{j}]_{q}}$
@@
The equivalence of (i) and (iii) implies (ii) because the cyclotomic polynomials are all monic. In the other direction, after dividing through by the leading coefficient which is also the greatest common divisor of all the coefficients by hypothesis, we can assume $f(q)$ is a monic polynomial with non-negative integer coefficients. If $f(q)$ also has all of its complex roots with modulus at most $1$, then the roots of $f$ are each either a root of unity or 00 by 3.1. Hence, (ii) implies (i).
@@
3.2. Rational products of $q$-integers

### v3

@@ 2.1 Cyclotomic polynomials
Also, $\Phi_{p}(q)=1+q+q^{2}+\cdots+q^{p-1}=[p]_{q}$ if and only if $p$ is prime. Note, $\Phi_{p^{k}}(1)=p$, otherwise if $n>1$ is not a prime power, then $\Phi_{n}(1)=1$. The number of complex roots and therefore the degree of $\Phi_{n}(q)$ is given by Euler’s totient function, $\varphi(n)$, so for $n\geqslant 3$ the degree of $\Phi_{n}(q)$ is even and bounded by $\sqrt{n/2}\leqslant\varphi(n)\leqslant n-1$. For $n\geqslant 2$, the constant term is $\Phi_{n}(0)=1$, and the coefficients are palindromic in the sense that $\Phi_{n}(q)=q^{\varphi(n)}\Phi_{n}(q^{-1})$.
2.2 Probabilistic generating functions and cumulants
We now briefly review the background from probability related to CGFs. See [Bil95] for more details or [BKS20a, §2] for a review aimed at a combinatorial audience.
The probability generating function of a discrete random variable $\mathcal{X}$ supported on $\mathbb{Z}_{\geqslant 0}$ is
@@ 2.2 Probabilistic generating functions and cumulants
If $\mathcal{X}$ is a nonnegative-integer-valued statistic on a set $S$ which is sampled uniformly, then the probability generating function of $\mathcal{X}$ as a random variable is the same as the ordinary generating function of $\mathcal{X}$, up to a scale factor:
@@
where $\mu_{d}\coloneqq\mathbb{E}[\mathcal{X}^{d}]$ is the $d$th moment. The central moments of $\mathcal{X}$ are $\alpha_{d}\coloneqq\mathbb{E}[(\mathcal{X}-\mu)^{d}]$, where $\mu=\mu_{1}$ is the mean and $\alpha_{2}=\sigma^{2}$ is the variance. The characteristic function of $\mathcal{X}$ is
@@
The normal distribution with mean $\mu$ and standard deviation $\sigma$, denoted $\mathcal{N}(\mu,\sigma)$ is central in this work. It is the unique distribution with cumumlants $\kappa_{1}=\mu,\kappa_{2}=\sigma^{2}$, and $\kappa_{d}=0$ for all positive integers $d\geqslant 3$.
@@
$f(q)=\frac{\prod_{k=1}^{m}[a_{k}]_{q}}{\prod_{k=1}^{m}[b_{k}]_{q}}=\sum_{k=0}^{n}c_{k}q^{k}\in\Phi^{+},$
so in particular each $c_{k}\in\mathbb{Z}_{\geqslant 0}$ and $n=\sum a_{k}-b_{k}\geqslant 0$. Let $\mathcal{X}$ be a discrete random variable with $\mathbb{P}[\mathcal{X}=k]=c_{k}/f(1)$. Then the $d$th cumulant of $\mathcal{X}$ is
$\kappa_{d}^{\mathcal{X}}=\frac{B_{d}}{d}\sum_{k=1}^{m}(a_{k}^{d}-b_{k}^{d}),$ (2.5)
@@
$\alpha_{d}=\sum_{\begin{subarray}{c}|\lambda|=d\\ \text{has all parts even}\end{subarray}}\frac{d!}{z_{\lambda}}\,\prod_{i=1}^{\ell(\lambda)}\frac{B_{\lambda_{i}}}{\lambda_{i}!}\left[\sum_{k=1}^{m}\left(a_{k}^{d}-b_{k}^{d}\right)\right],$ (2.6)
@@
$\mu_{d}=\sum_{\begin{subarray}{c}|\lambda|=d\\ \text{has all parts either}\\ \text{even or size $1$}\end{subarray}}\frac{d!}{z_{\lambda}}\,\prod_{i=1}^{\ell(\lambda)}\frac{B_{\lambda_{i}}}{\lambda_{i}!}\left[\sum_{k=1}^{m}\left(a_{k}^{d}-b_{k}^{d}\right)\right].$ (2.7)
Definition 10.
@@
Observe that $\mathcal{X}^{*}$ has mean $\mu^{*}=0$ and variance ${\sigma^{*}}^{2}=1$. The moments and central moments of $\mathcal{X}^{*}$ agree for $d\geqslant 2$ and are given by
@@
Similarly, the cumulants of $\mathcal{X}^{*}$ are given by $\kappa_{1}^{*}=0$, $\kappa_{2}^{*}=1$, and $\kappa_{d}^{*}=\kappa_{d}/\sigma^{d}$ for $d\geqslant 2$.
Definition 11.
@@
Theorem 12 (Lévy’s Continuity Theorem, [Bil95, Theorem 26.3]).
@@
Theorem 13 (Frechét–Shohat Theorem, [Bil95, Theorem 30.2]).
@@
By Theorem 12, we may test for asymptotic normality by checking if the standardized characteristic functions tend pointwise to the characteristic function of the standard normal. Likewise by Theorem 13 we may instead perform the check on the level of individual standardized moments, which is often referred to as the method of moments. By the polynomial relationship between moments and cumulants, we may further replace the moment condition (2.8) with the cumulant condition
$\lim_{n\to\infty}\kappa_{d}^{\mathcal{X}_{n}}=\kappa_{d}^{\mathcal{X}}.$ (2.9)
@@
Corollary 14.
A sequence $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ of real-valued random variables on finite sets is asymptotically normal if for all $d\geqslant 3$ we have
$\lim_{n\to\infty}\frac{\kappa^{\mathcal{X}_{n}}_{d}}{(\sigma^{\mathcal{X}_{n}})^{d}}=0.$ (2.10)
In fact, the converse of the Frechét–Shohat theorem holds for cyclotomic generating functions. See Theorem 40 below, which builds on [HZ15, Lem. 2.8]. Furthermore, we have the following simplified test for asymptotic normality due to Hwang and Zacharovas.
Theorem 15.
@@
$\lim_{n\to\infty}\frac{\kappa^{\mathcal{X}_{n}}_{4}}{(\sigma^{\mathcal{X}_{n}})^{4}}=0.$ (2.11)
2.3 Formal cumulants
@@
$\log f(e^{t})=\sum_{d=1}^{\infty}\kappa_{d}(f)\frac{t^{d}}{d!}$ (2.12)
@@
$\kappa_{d}(f)=\frac{B_{d}}{d}\left(\sum_{k=1}^{m}a_{k}^{d}-\sum_{k=1}^{m}b_{k}^{d}\right).$ (2.13)
@@
$\phi_{f}(t)\coloneqq f(e^{it})/f(1)=\frac{\prod_{k=1}^{m}[a_{k}]_{e^{it}}/a_{k}}{\prod_{k=1}^{m}[b_{k}]_{e^{it}}/b_{k}}=e^{-it\mu}\prod_{j=1}^{m}\frac{\operatorname{sinc}(a_{j}t/2)}{\operatorname{sinc}(b_{j}t/2)},$ (2.14)
@@
$\kappa_{d}(\Phi_{n}(q))=\frac{B_{d}}{d}\sum_{k\mid n}\mu(n/k)k^{d}=\frac{B_{d}}{d}J_{d}(n).$ (2.15)
@@
$\sum_{k\mid n}\mu(n/k)k^{d}=n^{d}\prod_{\begin{subarray}{c}\text{$p$ prime}\\ p\mid n\end{subarray}}\left(1-\frac{1}{p^{d}}\right).$ (2.16)
@@
Corollary 17.
[HZ15, Cor. 3.1] Let $f\in\mathbb{R}_{\geqslant 0}[q]$ be any nonzero polynomial such that all of its complex roots have modulus $1$. The corresponding odd cumulants $\kappa_{2d-1}(f)$ vanish after the first. The corresponding even cumulants $\kappa_{2d}(f)$ alternate in sign to make $(-1)^{d-1}\kappa_{2d}(f)\geqslant 0$.
2.4 Generalized uniform sum distributions
@@
Definition 18.
Let $\mathbf{t}=(t_{1},t_{2},\ldots)$ be a sequence of nonnegative real numbers. For $p\in\mathbb{R}_{\geqslant 1}$, the $p$-norm of $\mathbf{t}$ is $|\mathbf{t}|_{p}\coloneqq\left(\sum_{k=1}^{\infty}t_{k}^{p}\right)^{1/p}$. We also set $|\mathbf{t}|_{\infty}\coloneqq\sup_{k}t_{k}$.
The $p$-norm has many nice properties. It is well-known (e.g. [MV97, Ex. 7.3, p.58]) that if $1\leqslant p\leqslant q\leqslant\infty$, then $|\mathbf{t}|_{p}\geqslant|\mathbf{t}|_{q}$, and that if $|\mathbf{t}|_{p}<\infty$, then $\lim_{q\to\infty}|\mathbf{t}|_{q}=|\mathbf{t}|_{\infty}$. Thus, if $\mathbf{t}$ is weakly decreasing, $|\mathbf{t}|_{\infty}=\sup_{k}t_{k}=t_{1}$.
The sequence space with finite $p$-norm $\mathbf{\ell}_{p}\coloneqq\{\mathbf{t}=(t_{1},t_{2},\ldots)\in\mathbb{R}_{\geqslant 0}^{\mathbb{N}}:|\mathbf{t}|_{p}<\infty\}$ is commonly used in functional analysis and statistics. Here we define a related concept for analyzing sums of central continuous uniform random variables.
Definition 19.
@@
~$\widetilde{\ell}_{p}\coloneqq\{\mathbf{t}=(t_{1},t_{2},\ldots):t_{1}\geqslant t_{2}\geqslant\cdots\geqslant 0,|\mathbf{t}|_{p}<\infty\}.$
The elements of ~$\widetilde{\ell}_{p}$ may equivalently be thought of as the set of countable multisets of nonnegative real numbers with finite $p$-norm. Any finite multiset of nonnegative real numbers can be considered as an element of ~$\widetilde{\ell}_{p}$ with finite support by sorting the multiset and appending 00’s. We write ~$\tilde{\ell}_{\leqslant m}$ for the resulting collection of decreasing sequences with at most $m$ nonzero entries. The multisets in ~$\widetilde{\ell}_{p}$ are uniquely determined by their $p$-norms. In fact, any sequence of $p$-norm values injectively determines the multiset provided the sequence goes to infinity.
Definition 20.
@@
The uniform sum random variables have nice cumulant formulas which are similar to the CGF distributions. It was shown in [BS22, Lem. 3.11] that for $d\geqslant 2$ and $\mathbf{t}=(t_{1},t_{2},\ldots)\in\widetilde{\ell}_{2}$, we have
$\kappa_{d}^{\mathcal{S}_{\mathbf{t}}}=\frac{B_{d}}{d}\sum_{k=1}^{\infty}(t_{k})^{d}=\frac{B_{d}}{d}|\mathbf{t}|_{d}^{d}.$ (2.17)
Theorem 21.
@@
It is not known what all the possible limiting distributions of families of CGF polynomials are. By [BS22, Cor. 3.17], we know that all standardized uniform sum distributions $\mathcal{S}_{\mathbf{t}}^{*}$ do occur as limiting distributions coming from the hook length formulas for linear extensions of forests due to Björner and Wachs [BW89]. In fact, all standardized DUSTPAN distributions can occur as limits of CGF distributions. These are distributions of the form $\mathcal{S}_{\mathbf{t}}+\mathcal{N}(0,\sigma^{2})$, assuming the two random variables are independent, $\mathbf{t}\in\widetilde{\ell}_{2}$, and $\sigma^{2}\coloneqq 1-|\mathbf{t}|_{2}^{2}/12\in\mathbb{R}_{\geqslant 0}$.
3 Algebraic considerations
3.1 Equivalent characterizations of CGFs
@@
If $f(q)=\prod_{j=1}^{n}(q-z_{j})$ for $z_{1},\ldots,z_{n}\in\mathbb{C}$, define $f_{k}(q)=\prod_{j=1}^{n}(q-z_{j}^{k})$ for each positive integer $k$. The coefficients of $f_{k}(q)$ are elementary symmetric polynomials in the $z_{j}^{k}$ and each $|z_{j}^{k}|\leqslant 1$ by hypothesis, so the coefficients of $f_{k}(q)$ are bounded in modulus by binomial coefficients. These coefficients are also symmetric functions of roots of $f$, so they belong to the fixed field of the Galois group of $f$, namely the base field $\mathbb{Q}$. Since $f$ is monic, its roots $z_{1},\ldots,z_{n}$ are algebraic integers as are all sums of products of the $z_{i}$’s. Therefore, the coefficients of each $f_{k}(q)$ are also algebraic integers, and since the only algebraic integers in $\mathbb{Q}$ are the integers themselves, we observe $f_{k}(q)\in\mathbb{Z}[q]$. The list $f_{1},f_{2},\ldots$ must thus eventually repeat. We may as well suppose $f=f_{1}$ is repeated, so that taking $k$th powers for some $k>1$ permutes the $z_{i}$’s. Hence, taking $k$th powers $n!$ times implies $z_{i}=z_{i}^{kn!}$ for all $i$, and the result follows. ∎
@@
$\Phi_{n}(q)=\prod_{d\mid n}(q^{d}-1)^{\mu(n/d)}=\prod_{d\mid n}[d]_{q}^{\mu(n/d)}=\prod_{j=1}^{m}\frac{1-q^{a_{j}}}{1-q^{b_{j}}}=\prod_{j=1}^{m}\frac{[a_{j}]_{q}}{[b_{j}]_{q}}$ (3.1)
@@
The equivalence of (i) and (iii) implies (ii) because the cyclotomic polynomials are all monic. In the other direction, after dividing through by the leading coefficient which is also the greatest common divisor of all the coefficients by hypothesis, we can assume $f(q)$ is a monic polynomial with nonnegative integer coefficients. If $f(q)$ also has all of its complex roots with modulus at most $1$, then the roots of $f$ are each either a root of unity or 00 by Theorem 22. Hence, (ii) implies (i).
@@
3.2 Rational products of $q$,qintegers